OpenStudy (anonymous):

A circle of a d=22 miles what is the area of A to the nearest 10th use 3.14 to get the answer?

jhonyy9 (jhonyy9):

do you know formula of area of a circle ?

OpenStudy (anonymous):

No I don't understand math at all.

jhonyy9 (jhonyy9):

so A=pi*r^2 but here you know the diameter d=22 miles so from this how you can calcul the r ? so you know that the diameter is equal 2r than this mean that if you divide d/2 =r so than 22/2 =11 than r=11 so area will be pi*r^2 = 3,14*11^2 =3,14*121 = 379,94 miles squared

jhonyy9 (jhonyy9):

so this mean that the 3rd answer will be right sure ok ? do you understand it ?

jhonyy9 (jhonyy9):

@katebored are you here ?

OpenStudy (anonymous):

Yes trying to understand. But I don't lol

jhonyy9 (jhonyy9):

what you dont understand here ?

OpenStudy (anonymous):

You were right. Thank you

jhonyy9 (jhonyy9):

was my pleasure bye